OpenStudy (anonymous):

Given that f(x) = x2 + 2x + 3 and g(x) = quantity of x plus four, over three, solve for f(g(x)) when x = 2. 2 5 11 33

OpenStudy (anonymous):

IS IT

OpenStudy (aum):

f(x) = x^2 + 2x + 3 g(x) = (x+4)/3 f(g(x)) when x = 2 is f(g(2)) Evaluate the innermost function first. Find g(2). Then take that value and evaluate f(x) at that value.

OpenStudy (anonymous):

11

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@mitchal

OpenStudy (aum):

Yes, 11.